OpenStudy (anonymous):

I fan and medal!!! :) find the slope of a line that passes through (-2 -3) and (1 1).

OpenStudy (mathmale):

Look for the formula for "slope of a straight line." I'd bet you've seen and used this before. Find the slope of the line thru the given points, showing all work, please.

OpenStudy (anonymous):

m=y2-y1 over x2-x1 ? I'm not sure how to set it up like which numbers goes where..

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

You are given two points: (-2 -3) and (1 1) Either one could be regard as Point #1 and then the other regarded as Point #2. If Point #1 is (-2,-3), then x1 = -2 and y1=-3. Tray again. Find the slope of the line connecting these two points.

OpenStudy (anonymous):

-3-(-2) = -1 so would the slope be -1?

OpenStudy (anonymous):

oh wait but then there is (1,1) okay i will do that one aswell

OpenStudy (anonymous):

1-1=0 so -1?? @mathmale

OpenStudy (mathmale):

Please show your work. Use the formula for the slope: \[m=\frac{ y-y _{1} }{ x-x _{1} }\] First, identify which point is Point#1 and which is Point #2.

OpenStudy (mathmale):

You are correct here: "m=y2-y1 over x2-x1 ? I'm not sure how to set it up like which numbers goes where.."